OpenStudy (anonymous):

(7+sq root of 2)(8- 3 sq root 2)

OpenStudy (anonymous):

simplify is the question ??

OpenStudy (anonymous):

square root of 2 is =1.414

OpenStudy (anonymous):

not as a decimal, simply by like multiplying those two binomials. i keep getting the wrong answer

OpenStudy (mertsj):

\[(7+\sqrt{2})(8-3\sqrt{2})\]

OpenStudy (mertsj):

Use FOIL...multiply the two binomials.

OpenStudy (anonymous):

kk @Mertsj ... gave u hint ....

OpenStudy (anonymous):

i used foil and got the wrong answer @Mertsj

OpenStudy (anonymous):

whats the ans ....???

OpenStudy (anonymous):

i don't know what the right answer is i just know that what i got is wrong

OpenStudy (anonymous):

i got 56-10 sq root 4

OpenStudy (anonymous):

is the ans \[50-13\sqrt{2}\] i got this ....

OpenStudy (anonymous):

maybe! how did you get that?

OpenStudy (anonymous):

\[(7+\sqrt{2})(8-3\sqrt{2}) = 56-21\sqrt{2}+8\sqrt{2}-6 = 50 -13\sqrt{2}\] ....ok

OpenStudy (anonymous):

kkk na @heyitsmelalala

OpenStudy (anonymous):

Good job @shkrina

OpenStudy (anonymous):

thanbku ....

OpenStudy (anonymous):

You are always welcome